OpenStudy (anonymous):

Choose the equation below whose axis of symmetry is x = 0. y = x2 + 2x y = x2 − 16x + 58 y = x2 + 2 y = x2 − 4x + 2

OpenStudy (anonymous):

x = -b/2a = 0 <=> b = 0 Can you choose now?

OpenStudy (anonymous):

idk

OpenStudy (anonymous):

what's b?

OpenStudy (anonymous):

y = x2 − 16x + 58

OpenStudy (anonymous):

y = x² + 2x In this equation b = 2

OpenStudy (anonymous):

Choose the option with b = 0

OpenStudy (anonymous):

is the answer b ?

OpenStudy (anonymous):

y = x²− 16x + 58 Nope, b = -16

OpenStudy (anonymous):

c?

OpenStudy (anonymous):

y = x² + 2 Yes, no b = 0 here

OpenStudy (anonymous):

Yes, no b => b = 0

OpenStudy (anonymous):

Its c ? youre confusing me ;o llol